# Knowledge Graph Based Agent for Complex, Knowledge-Intensive QA in Medicine

### v1

@@ Abstract
Biomedical knowledge is uniquely complex and structured, requiring distinct reasoning strategies compared to other scientific disciplines like physics or chemistry. Biomedical scientists do not rely on a single approach to reasoning; instead, they use various strategies, including rule-based, prototype-based, and case-based reasoning. This diversity calls for flexible approaches that accommodate multiple reasoning strategies while leveraging in-domain knowledge. We introduce KGARevion, a knowledge graph (KG) based agent designed to address the complexity of knowledge-intensive medical queries. Upon receiving a query, KGARevion generates relevant triplets by using the knowledge base of the LLM. These triplets are then verified against a grounded KG to filter out erroneous information and ensure that only accurate, relevant data contribute to the final answer. Unlike RAG-based models, this multi-step process ensures robustness in reasoning while adapting to different models of medical reasoning. Evaluations on four gold-standard medical QA datasets show that KGARevion improves accuracy by over 5.2%, outperforming 15 models in handling complex medical questions. To test its capabilities, we curated three new medical QA datasets with varying levels of semantic complexity, where KGARevion achieved a 10.4% improvement in accuracy.
@@ 1 Introduction
Medical reasoning involves making diagnostic and therapeutic decisions while also understanding the pathology of diseases (Patel et al., 2005). Unlike many other scientific domains, medical reasoning often relies on vertical reasoning, using analogy more heavily (Patel et al., 2005). For instance, in biomedical research, an organism such as Drosophila is used as an exemplar to model a disease mechanism, which is then applied by analogy to other organisms, including humans. In clinical practice, the patient serves as an exemplar, with generalizations drawn from many overlapping disease models and similar patient populations (Charles et al., 1997; Menche et al., 2015). In contrast, fields like physics and chemistry tend to be horizontally organized, where general principles are applied to specific cases (Blois, 1988). This distinction highlights the unique challenges that medical reasoning poses for question-answering (QA) models.
@@
While large language models (LLMs) (OpenAI, 2024; Dubey et al., 2024; Gao et al., 2024) have demonstrated strong general capabilities, their responses to medical questions often suffer from incorrect retrieval, missing key information, and misalignment with current scientific and medical knowledge. Additionally, they can struggle to provide contextually relevant answers that account for specific local contexts, such as patient demographics or geography, as well as specific areas of biology (Harris, 2023). A major issue lies in these models’ inability to systematically integrate different types of evidence. Specifically, they have difficulty combining scientific factual (structured, codified) knowledge derived from formal, rigorous research with tacit (noncodified) knowledge—expertise and lessons learned—which is crucial for contextualizing and interpreting scientific evidence in relation to the specific modifying factors of a given medical question (Harris, 2023).
LLM-powered QA models often lack such multi-source and grounded knowledge necessary for medical reasoning, which requires understanding the nuanced and specialized nature of medical concepts. Additionally, LLMs trained on general knowledge may struggle to solve medical problems that demand specialized in-domain knowledge. This shortcoming arises from their inability to discern subtle, granular differences that are critical in medical contexts. As a result, LLMs face challenges in complex medical reasoning because such reasoning requires both: 1) simultaneous consideration of dependencies across multiple medical concepts within an input question, and 2) precise, local in-domain knowledge of semantically similar concepts that can carry different medical meanings, as we demonstrate in Fig. 1.
The prevailing strategy to address these challenges is the use of information retrieval techniques, such as retrieval-augmented generation (RAG), which follows a Retrieve-then-Answer paradigm (Xiong et al., 2024b; Shi et al., 2024). Although these methods can provide multi-source knowledge from external databases (Fan et al., 2024), the accuracy of the generated answers depends heavily on the quality of the retrieved information, making them vulnerable to potential errors (Karpukhin et al., 2020). Data repositories and knowledge bases these models draw from contain incomplete or incorrect information, leading to inaccurate retrieval Adlakha et al. (2024); Thakur et al. (2023). Further, many RAG-based methods lack post-retrieval verification mechanisms to validate that retrieved information is factually correct and does not miss key information (Zhao et al., 2023). Knowledge graphs (KGs) of medical concepts have been widely adopted as a grounded knowledge base to provide precise and specialized in-domain knowledge for medical QA models (Qi et al., 2024; Murali et al., 2023; Chandak et al., 2023; Jiang et al., ). While KGs can enhance the performance of these models, they are often incomplete. Consequently, approaches that retrieve medical concepts from a KG based solely on the presence of direct associations (edges) between concepts are insufficient. For instance, concepts representing two proteins with distinct biological roles may not be directly connected in the KG, even though these proteins share similar biological representations (Menche et al., 2015). To advance LLM-powered models for knowledge-intensive medical QA, it is essential to develop models that can (1) consider complex associations between several medical concepts at the same time, (2) systematically integrate multi-source knowledge, and (3) effectively verify and ground the retrieved information to ensure contextual relevance and accuracy.
@@
Figure 1: a) Performance of LLMs on three new datasets introduced in this paper with questions of varying difficulty. b) Sample questions from new MedDDx-Basic, MedDDx-Intermediate, and MedDDx-Expert datasets.
@@ 2 Related work
LLM-based reasoning. General-purpose LLMs (GPT (OpenAI, 2024), LLaMA family (Dubey et al., 2024; Touvron et al., 2023), Mistral (Jiang et al., 2023)), and LLMs fine-tuned on biomedical data (BioMedLM (Venigalla et al., 2022), Codex (Liévin et al., 2024), MedAlpaca (Han et al., 2023), Med-PaLM (Singhal et al., 2023), PMC-LLaMA (Wu et al., 2024a)) are used for medical reasoning by leveraging their vast embedded knowledge. Other models utilize the open-ended reasoning capabilities of LLMs to break down queries into sub-tasks, arriving at the final answer step by step, such as Chain-of-Though (CoT) (Wei et al., 2024), CODEX COT (Gramopadhye et al., 2024). However, these methods often struggle with knowledge-intensive medical queries requiring multi-sources an specific knowledge.
RAG-based models. Self-RAG (Asai et al., 2024) is a pioneering framework that enhances LLM performance through retrieval and self-reflection. LLM-AMT (Wang et al., 2023b) improves medical question answering by integrating authoritative medical textbooks into large language models with specialized knowledge retrieval and self-refinement techniques. Adaptive-RAG (Jeong et al., 2024) introduces a dynamic RAG framework that adapts retrieval strategies based on question complexity. However, its accuracy is constrained by the quality of retrieved knowledge (Zhang et al., 2024).
KG-based models. Before the rise of LLMs, several models, such as QAGNN (Yasunaga et al., 2021), JointLK (Sun et al., 2022), and Dragon (Yasunaga et al., 2022), were developed to tackle medical queries solely using KGs in an end-to-end manner. However, these methods cannot be easily applied to questions involving unseen nodes or incomplete knowledge within the graphs. In addition, KGs, with their structured and reliable information, have driven research toward RAG models based on graph data, motivating models like GraphRAG (Edge et al., 2024), KG-RAG (Soman et al., 2023), and MedGraphRAG (Wu et al., 2024b). To improve retrieval accuracy, KG-Rank (Yang et al., 2024) is introduced to rank retrieved triplets and filter out irrelevant knowledge. Additionally, GenGround (Shi et al., 2024) uses a Generate-then-Ground pipeline that grounds answers by prompting LLMs to validate retrieved knowledge. However, all these approaches rely heavily on semantic dependencies, overlooking the rich structural information within KGs.
3 Approach
@@
To address this problem, we propose developing an LLM-powered agent framework (Wu et al., 2023; Li et al., 2023) that leverages various actions (Schick et al., 2023; Shen et al., 2023; Nakano et al., 2021) to collaboratively perform complex tasks (Tang et al., 2023; Bran et al., 2023; Boiko et al., 2023). Fig. 2 shows an overview of KGARevion, which comprises four key actions, including Generate (§3.1), Review (§3.2), Revise (§3.3), and Answer (§3.3) actions. The Generate action is responsible for generating triplets related to the input question. The Review action then assesses the correctness of each generated triplet, while the Revise action corrects any triplet identified as being incorrect. Finally, the Answer action outputs the final answer based on the triplets identified as correct by the Review action.
Figure 2: a) The overview of KGARevion. b) The architecture of fine-tuning stage in the Review action, where embeddings get from KGs are structural embeddings, while concept embeddings from LLMs.
@@ 3.1 Generate Action
Depending on the content of each answer candidate $a\in C$, the input questions can be broadly categorized into two types: choice-aware and non-choice-aware. The answer candidates in the choice-aware group have specific contents, whereas the ones in the non-choice-aware group only contain yes-or-no options. These different types of questions require distinct reasoning processes: choice-aware questions involve analyzing the content of each answer candidate, while non-choice-aware questions only require focusing on the question stem.
To handle this, the Generate action is designed to prompt the LLM (Ouyang et al., 2022; Wang et al., 2023a) to follow different procedures for generating relevant triplets according to the type of input question.
@@
where the LLM is prompted to extracts triplets from the medical concepts involved in input question stem $q$.
@@ 3.2 Review Action
To enable LLMs to accurately judge the correctness of generated triplets, beyond relying solely on semantic dependencies inferred by LLMs (Shinn et al., 2023), the Review action also leverages the relationships among various medical concepts contained in KGs. This is achieved by fine-tuning the LLM on a KG completion task, explicitly integrating entity structural embeddings learned from KGs into the LLM. Once fine-tuned, the Review action utilizes the model to assess the correctness of generated triplets, as shown in Fig. 2.
To enable the LLM to capture the structural information embedded in KGs, we employ TransE (Bordes et al., 2013) to learn structural embeddings for both entities and relations in $G$. These obtained embeddings are fixed when fine-tuning the LLM on knowledge completion tasks.
Fine-tuning stage. Given a triplet $(h,r,t)\in G$ and its pre-trained embedding $\textbf{e}_{h}\in\mathbb{R}^{d}$, $\textbf{e}_{r}\in\mathbb{R}^{d}$, and $\textbf{e}_{t}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedding, the aim of this stage is to learn a function $f(\cdot)$ that makes the LLM capable of determining if a given triplet is True or False by considering its structural embeddings in the KG, as follows:
$b=f(P(\cdot),(h,r,t),\mathbf{e}_{h},\mathbf{e}_{r},\mathbf{e}_{t}),~{}b\in\{\textrm{True},\textrm{False}\}$ (2)
where $b$ is a bool value (True or False) and $P(\cdot)$ denotes the LLM.
However, it is complex for LLMs to directly understand structural embeddings given that the embeddings in LLMs are obtained based on their vocabularies (Radford et al., 2019). To make sure that LLMs can understand structural embeddings, we generate a description for each triplet based on the relationship between the two entities. Then, we tokenize the generated descriptions (Touvron et al., 2023) and obtain the corresponding token embeddings. The token embeddings are aligned with the pre-trained structural embeddings of the triplets to produce new embeddings for each input triplet.
In detail, given the input triplet $(h,r,t)$, we denote its description as $D(r)$, where $D$ represents the description dictionary with the key as relation and the value as a pre-defined description template, as shown in Appendix 8. Assuming the token embedding of $D(r)$ obtained from LLM $P(\cdot)$ is $\mathbf{X}\in\mathbb{R}^{|l|\times d_{p}}$, where $|l|$ denotes the number of max tokens and $d_{p}$ denotes the dimension of the embeddings in $P(\cdot)$, we first adopt a linear layer ($g(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{p}}$) to map the dimension of pre-trained embeddings of entities and relations to the same dimension of that of token embeddings. Then, we concatenate them to create the triplet embedding matrix $\mathbf{V}=[g(\mathbf{e}_{h});g(\mathbf{e}_{r});g(\mathbf{e})_{t})]\in\mathbb{R}^{3\times d_{p}}$. Afterwards, we adopt an attention block (Vaswani, 2017), followed by a two-layer feedforward neural network (denoted as FFN in Fig. 2b) to obtain the aligned triplet embedding matrix $\mathbf{Z}\in\mathbb{R}^{3\times d_{p}}$, as follows:
@@
where $\sigma(\cdot)$ denotes the $\mathrm{Softmax}$ function, $\varphi(\cdot)$ denotes the layer normalization function,$\mathbf{W}_{1}\in\mathbb{R}^{d_{p}\times d_{h}}$ and $\mathbf{W_{2}}\in\mathbb{R}^{d_{h}\times d_{p}}$ denotes the trainable parameters in the two-layer FFN, and $d_{h}$ denotes the dimension of hidden layer in the FFN.
The obtained aligned triplet embedding matrix is input into the LLM in the form of three prefix tokens together with an instruction $s$ to fine-tune the LLM to execute the knowledge graph completion task. Then, the Eq. 2 could be rewritten by:
$b=f(P(\cdot),D(r),g(\mathbf{e}_{h}),g(\mathbf{e}_{r}),g(\mathbf{e}_{t}))=f(P(\cdot),\mathbf{X},\mathbf{V})=f(P(\mathbf{Z},s))$ (5)
During the fine-tuning stage, the pre-trained entity and relation embeddings are frozen, and LoRA (Hu et al., 2022) is adopted to fine-tune the LLM. The trainable parameters are optimized using the next token prediction loss (Radford, 2018).
Inference stage. After fine-tuning, the model is used to evaluate the triplets in $T$ derived from the Generate action (3.1). In detail, we first adopt UMLS code (Bodenreider, 2004) to map entities in each triplet $(h,r,t)\in T$ to KGs to get the corresponding pre-trained entity and relation embeddings, denoted as $\mathbf{e}_{h},\mathbf{e}_{r},\mathbf{e}_{t}$. Then, the triplets are input into the fine-tuned LLM to determine if they are correct or not by Eq. 5.
However, not all entities in generated triplet $(h,r,t)\in T$ can be mapped to those in KGs. To address this, the Review action applies a soft constraint rule to distinguish whether the generated triplet is factually incorrect or the result of incomplete knowledge in KGs, as follows:
  • Factually Wrong: if we can map $h$ and $t$ to entities in KGs and $b=\mathrm{False}$, then the triplet $(h,r,t)$ is factually wrong and is removed from $T$.

### v2

# KGARevion: An AI Agent for Knowledge-Intensive Biomedical QA

@@ Abstract
Biomedical reasoning integrates structured, codified knowledge with tacit, experience-driven insights. Depending on the context, quantity, and nature of available evidence, researchers and clinicians use diverse strategies, including rule-based, prototype-based, and case-based reasoning. Effective medical AI models must handle this complexity while ensuring reliability and adaptability. We introduce KGARevion, a knowledge graph-based agent that answers knowledge-intensive questions. Upon receiving a query, KGARevion generates relevant triplets by leveraging the latent knowledge embedded in a large language model. It then verifies these triplets against a grounded knowledge graph, filtering out errors and retaining only accurate, contextually relevant information for the final answer. This multi-step process strengthens reasoning, adapts to different models of medical inference, and outperforms retrieval-augmented generation-based approaches that lack effective verification mechanisms. Evaluations on medical QA benchmarks show that KGARevion improves accuracy by over 5.2% over 15 models in handling complex medical queries. To further assess its effectiveness, we curated three new medical QA datasets with varying levels of semantic complexity, where KGARevion improved accuracy by 10.4%. The agent integrates with different LLMs and biomedical knowledge graphs for broad applicability across knowledge-intensive tasks. We evaluated KGARevion on AfriMed-QA, a newly introduced dataset focused on African healthcare, demonstrating its strong zero-shot generalization to underrepresented medical contexts.
@@ 1 Introduction
Biomedical reasoning requires integrating diagnostic and therapeutic decision-making with an understanding of the biology and chemistry of the disease of drugs (Patel et al., 2005). Large language models (LLMs)(OpenAI, 2024; Dubey et al., 2024; Gao et al., 2024) demonstrate strong general capabilities, but their responses to medical questions often suffer from incorrect retrieval, omission of key information, and misalignment with current scientific and clinical knowledge. These models can struggle to generate contextually relevant answers that account for local factors, such as patient demographics, geographic variations, and specialized biomedical domains(Harris, 2023). A limitation arises from their inability to integrate multiple types of evidence by combining structured, codified scientific knowledge with tacit, noncodified expertise—clinical intuition, case-based experience, and learned heuristics, which are essential for contextualizing scientific evidence within real-world medical decision-making (Harris, 2023).
LLM-powered QA models often lack the multi-source and grounded knowledge required for effective medical reasoning. The answer to complex medical questions requires an understanding of the intricate and highly specialized nature of biomedical concepts. However, LLMs trained in general-purpose data struggle to solve medical problems that require specialized knowledge in the domain. This challenge stems from their inability to differentiate subtle, domain-specific nuances of medical significance. As a result, LLMs often do not reason effectively in complex scenarios, where successful inference requires recognizing and reasoning about dependencies between multiple interrelated medical concepts within a single question and interpreting highly similar yet semantically distinct biomedical entities with precision, as illustrated in Fig. 1.
To address these limitations, researchers have turned to retrieval-augmented generation (RAG), which follows a retrieve-then-answer paradigm (Shi et al., 2024). These methods enrich LLMs with external knowledge sources, allowing them to retrieve information from biomedical databases and structured repositories (Fan et al., 2024). However, the accuracy of RAG-based answers is heavily dependent on the quality of the retrieved content, making them prone to errors (Karpukhin et al., 2020). Many biomedical knowledge bases contain incomplete, outdated, or incorrect information, leading to unreliable retrieval Adlakha et al. (2024); Thakur et al. (2023). Furthermore, approaches based on RAG lack post-retrieval verification mechanisms, leaving them unable to assess whether retrieved information is factually accurate, contextually relevant, or if it omits essential information (Zhao et al., 2023).
@@
Figure 1: a) Performance of existing LLMs on three new datasets (MedDDx-Basic, MedDDx-Intermediate, MedDDx-Expert) introduced in this paper with questions of varying difficulty. b) Sample questions from the new datasets.
Knowledge graphs (KG) of medical concepts serve as grounded knowledge bases that provide precise and specialized in-domain knowledge for medical QA models (Qi et al., 2024; Murali et al., 2023; Chandak et al., 2023). Although KGs improve these models, they often contain gaps, limiting their ability to capture biomedical relationships. Approaches that retrieve medical concepts based solely on direct associations (edges) in a KG fail to account for implicit or complex relationships. For example, two proteins with different biological functions may lack a direct connection in a KG, even though they share biological similarity at the molecular level (Menche et al., 2015). Advancing LLM-powered models for knowledge-intensive medical QA requires systems that simultaneously capture complex associations among multiple medical concepts, integrate multi-source knowledge systematically, and verify retrieved information to ensure contextual accuracy and relevance.
Present work. We introduce KGARevion, a KG-based LLM agent designed for complex biomedical QA that integrates the non-codified knowledge of LLMs with the structured, codified knowledge found in KGs. As illustrated in Fig. 2, KGARevion executes four key actions to ensure accurate and context-aware biomedical reasoning. First, it prompts the LLM to generate relevant triplets based on the input question. To effectively leverage structured KG data, KGARevion fine-tunes the LLM on a KG completion task, incorporating pre-trained structural embeddings of triplets as prefix tokens. The fine-tuned model then evaluates the correctness of the generated triplets. Next, KGARevion performs a ‘Revise’ action to correct erroneous triplets, refining the knowledge base before selecting the final answer. Given the complexity of medical reasoning, KGARevion adaptively chooses the most appropriate reasoning approach for each question, allowing for nuanced and context-aware QA. This flexibility enables KGARevion to handle both multi-choice and open-ended QA tasks.
@@ 2 Related Work
LLM-based reasoning. General-purpose LLMs (GPT (OpenAI, 2024), LLaMA family (Dubey et al., 2024; Touvron et al., 2023), Mistral (Jiang et al., 2023)) and LLMs fine-tuned on biomedical data (BioMedLM (Venigalla et al., 2022), Codex (Liévin et al., 2024), MedAlpaca (Han et al., 2023), Med-PaLM (Singhal et al., 2023), PMC-LLaMA (Wu et al., 2024a)) leverage their vast embedded knowledge to perform medical reasoning. Some models enhance reasoning by decomposing complex queries into sub-tasks, solving them step by step using structured prompts, as seen in Chain-of-Thought (CoT) (Wei et al., 2024) and CODEX CoT (Gramopadhye et al., 2024). However, these methods struggle with knowledge-intensive medical queries that require multi-source domain-specific knowledge, leading to gaps in accuracy and completeness.
RAG-based models. Self-RAG (Asai et al., 2024) is a framework that enhances LLM performance through retrieval and self-reflection. LLM-AMT (Wang et al., 2023b) improves medical question answering by integrating authoritative medical textbooks into LLMs with specialized knowledge retrieval and self-refinement techniques. Adaptive-RAG (Jeong et al., 2024) introduces a dynamic RAG framework that adapts retrieval strategies based on the complexity of the questions. However, its performance is restricted by the quality of the knowledge retrieved (Zhang et al., 2024).
KG-based models. Models, such as QAGNN (Yasunaga et al., 2021), JointLK (Sun et al., 2022), and Dragon (Yasunaga et al., 2022), handle medical questions solely using KGs in an end-to-end manner. However, these methods cannot be easily applied to questions involving unseen nodes or incomplete knowledge within the graphs. In addition, structured KGs have driven research toward graph-based RAG models, motivating models such as GraphRAG (Edge et al., 2024), KG-RAG (Soman et al., 2023), and MedGraphRAG (Wu et al., 2024b). KG-Rank (Yang et al., 2024) ranks the retrieved triplets and filters out irrelevant knowledge to improve the accuracy of the search. Additionally, GenGround (Shi et al., 2024) uses a Generate-then-Ground pipeline that grounds answers by prompting LLMs to validate retrieved knowledge. However, these approaches rely heavily on semantic dependencies, overlooking the rich structural information within the KGs.
3 KGARevion Agent
@@
KGARevion is an LLM-powered agent (Wu et al., 2023; Li et al., 2023) that defines agentic actions (Schick et al., 2023; Shen et al., 2023; Nakano et al., 2021) to collaboratively perform complex tasks (Tang et al., 2023; Bran et al., 2023; Boiko et al., 2023). Fig. 2 provides an overview of KGARevion, which operates through four key actions: Generate (§3.1; Generates triplets relevant to the input question), Review (§3.2; Assesses the correctness of each generated triplet), Revise (§3.3; Corrects any triplet identified as incorrect), and Answer (§3.3; Produces the final answer based on the triplets verified by the Review action). This structured reasoning process allows KGARevion to integrate both LLM-generated knowledge and structured KG-based validation, improving accuracy and robustness in knowledge-intensive medical QA.
Figure 2: a) Overview of KGARevion agent. b) Overview of fine-tuning in the Review action.
@@ 3.1 Generate Action
Depending on the content of the answer candidate, the input questions can be broadly categorized into two types: choice-aware and non-choice-aware. The answer candidates in the choice-aware group have specific contents, whereas the ones in the non-choice-aware group only contain yes-or-no options (as shown in Appendix Table 4). These different types of questions require distinct reasoning processes: choice-aware questions involve analyzing the content of each answer candidate, while non-choice-aware questions only require focusing on the question stem.
To handle this, this action is designed to prompt the LLM (Ouyang et al., 2022; Wang et al., 2023a) to follow different procedures for generating relevant triplets according to the input question type.
@@
where $a_{i}$ denotes the candidate answer in $C$, and the LLM $P(\cdot)$ is prompted to extracts triplets from the medical concepts involved in its input.
@@ 3.2 Review Action
To enable LLMs to accurately judge the correctness of generated triplets, beyond relying solely on semantic dependencies inferred by LLMs  (Shinn et al., 2023), the Review action also leverages the connections and relationships among various medical concepts contained in KGs. This is achieved by fine-tuning the LLM on a KG completion task, explicitly integrating entity embeddings learned from KGs into the LLM. Then the Review action is performed by the fine-tuned LLM to assess the correctness of triplets generated by the Generate action, as shown in Fig. 2.
3.2.1 Fine-tuning Stage
Generating KG Embeddings and Triplet Descriptions. We use the well-known KG representation learning method, TransE (Bordes et al., 2013), to learn structural embeddings for both entities and relations in $G$. For a triplet $(h,r,t)\in G$, the learned pre-trained embeddings are denoted as $\textbf{e}_{h}\in\mathbb{R}^{d}$, $\textbf{e}_{r}\in\mathbb{R}^{d}$, and $\textbf{e}_{t}\in\mathbb{R}^{d}$, where $d$ represents the embedding dimension. These embeddings are kept fixed during LLM fine-tuning. In addition, we instruct the LLM to generate a description template for each relation $r\in G$. We store these descriptions in a dictionary $D(\cdot)$, which can be found in Appendix Table 8.
Aligning Embedding. Since the embeddings in LLMs are based on token vocabularies (Radford et al., 2019), LLMs cannot directly interpret pre-trained structural embeddings, as they lack semantic meaning. To make use of the pre-trained structural embeddings, we align them with the corresponding descriptions to generate new embeddings for the input triplet.
Specifically, given the description $D(r)$ for the input triplet $(h,r,t)$, we denote the embedding of $D(r)$ obtained from the LLM as $\mathbf{X}\in\mathbb{R}^{|l|\times d_{L}}$, where $|l|$ is the maximum number of tokens and $d_{L}$ is the embedding dimension in the LLM. Next, we concatenate the embeddings of the head entity, the relation, and the tail entity, denoted as $\mathbf{V}=[g(\mathbf{e}_{h});g(\mathbf{e}_{r});g(\mathbf{e}_{t})]\in\mathbb{R}^{3\times d_{L}}$, where $g(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{L}}$. We then apply an attention block (Vaswani, 2017), followed by a two-layer feed-forward network (FFN, as shown in Fig. 2b) to obtain the aligned triplet embedding matrix $\mathbf{Z}\in\mathbb{R}^{3\times d_{L}}$ as follows:
@@ 3.2.1 Fine-tuning Stage
where $\sigma(\cdot)$ is the Softmax function, $\varphi(\cdot)$ represents layer normalization, $\mathbf{W}_{1}\in\mathbb{R}^{d_{L}\times d_{h}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{h}\times d_{L}}$ are trainable parameters in the FFN, and $d_{h}$ is the dimension of the hidden layer in the FFN.
Fine-tuning LLM. After obtaining the aligned embedding $\mathbf{Z}$, we add it to the beginning of the instruction and fine-tune the LLM using LoRA (Hu et al., 2022) with the next-token prediction loss (Radford, 2018). The instruction is: ’Given a triple from a knowledge graph. Each triple consists of a head entity, a relation, and a tail entity. Please determine the correctness of the triple and response with True or False.’ The output should be either True or False.
3.2.2 Inference Stage
The fine-tuned LLM is then integrated to the Review action to check the accuracy of each triplet in $T$, which was generated by the Generate action (3.1). Specifically, we first use UMLS codes (Bodenreider, 2004) to map entities in the KG and obtain pre-trained structural embeddings for the head entity, relation, and tail entity, respectively. These embeddings, along with their descriptions and instructions, are fed into the fine-tuned LLM to determine whether the triplet is correct or not.
However, not all entities in the generated triplet $(h,r,t)\in T$ can be mapped to entities in KGs. To address this, the Review action applies a soft constraint rule to distinguish whether the generated triplet is factually wrong or the result of incomplete knowledge in KGs, as follows:
  • Factually Wrong: if we can map $h$ and $t$ to entities in KGs and the output of fine-tuned LLM is False, then the triplet $(h,r,t)$ is factually wrong and is removed from $T$.